Roger's age is two times the age of his sister. Last year his age was three times the age of his sister. How old is Roger?
Answer by Shana-D77(132)

This year, Roger is R and Sister is S
Last year, Roger is R-1 and Sister is S-1
TODAY:
R = 2S
LAST YEAR:
R - 1 = 3(S - 1)
R - 1 = 3S - 3
Sub in 2S for R (because R = 2S)
2S - 1 = 3s - 3
S = 2. So sister is 2 this year, Roger is 4.
Last year, Sister was 1 and Roger was "three times her age" or 3.